# Aspects of the commuting graph [^1] To the memory of Richard Parker

### v3

@@ 3 The Recognition Problem
    For the general case, we obtain a quasipolynomial-time algorithm. The algorithm is essentially based on a simple $2^{O(\log^{3}n)}$ time algorithm that enumerates the multiplication tables of all distinct groups of order $n$ (where by distinct we mean mutually non-isomorphic groups of order $n$). This enumeration algorithm combined with Babai’s quasipolynomial time algorithm for graph isomorphism [4] yields a simple $2^{O(\log^{3}n)}$ time algorithm to check if a given $n$ vertex graph $X$ is the commuting graph of an $n$-element group: for each $n$-element group $G$ that is output by the enumeration algorithm we check if $\Gamma(G)$ is isomorphic to $X$ by using Babai’s isomorphism test [4].
@@ 4 Basic properties
A vertex subset $S$ is a maximal clique of $X$ iff $S$ is a maximal abelian subgroup of $G$.
Proof 1
Suppose $H$ is a maximal abelian subgroup of $G$. Clearly, $H$ is a clique in $X$. If $x\notin H$ is adjacent to all of $H$ then $x$ commutes with all of $H$ implying that $\langle H\cup\{x\}\rangle$ is an abelian subgroup of $G$ strictly larger than $H$. Hence the clique induced by $H$ is maximal. Conversely, by a similar argument, if $S$ is a maximal clique in $X$ then $S$ is a maximal abelian subgroup of $G$$\qed$
@@ 4.1 The vertex degrees of $\Gamma(G)$ and conjugacy classes of $G$
An $n$-vertex graph $X$ with fewer than $n(f(n)-1/2)$ edges cannot be the commuting graph of a group, where $f(n)$ is the minimum number of conjugacy classes that an $n$ element group can have.
At the other extreme, we can rule out very dense incomplete graphs by the $5/8$-theorem [41] for finite groups. More precisely, for any finite nonabelian group $G$ with $n$ elements:
@@ 4.2 The commuting graph and maximal abelian subgroups
For a finite group $G$, the number of maximal abelian subgroups is bounded by (${|G|}\choose{\log|G|}$ because every subgroup of $G$ has a generating set of size bounded by $\log|G|$. Thus, the hypergraph of maximal abelian subgroups has size at most $n^{\log n}$ for $n$ element groups.
@@ 5 Commuting graphs of product groups
[42, Chapter 4, Page 35] Let $X=(V,E)$ and $X^{\prime}=(V^{\prime},E^{\prime})$ be simple undirected graphs. Their strong product, denoted $X\boxtimes X^{\prime}$, is a simple undirected graph with the cartesian product $V\times V^{\prime}$ as its vertex set and edges defined as follows: distinct pairs $(u,u^{\prime})$ and $(v,v^{\prime})$ are adjacent if and only if one of the following holds:
@@
The strong product is both commutative and associative [42, Chapter 4], in the sense that $X\boxtimes Y\cong Y\boxtimes X$ and $(X\boxtimes Y)\boxtimes Z)\cong X\boxtimes(Y\boxtimes Z)$.
By a nontrivial graph we mean a graph with more than one vertex. Given a graph $X$, a factorization $X\cong Y\boxtimes Z$ is nontrivial if both $Y$ and $Z$ are nontrivial. A graph $X$ is said to be a prime graph if it does not have a nontrivial factorization. (We warn the reader that this term is also used for graphs defined by arithmetic parameters of a finite group; see for example [53].)
It turns out that any connected graph $X$ has a unique factorization (up to isomorphism) into a strong product of prime graphs [42, Theorem 7.14, Pg 85]. Moreover, this prime factorization can be computed in polynomial time [34]. As a consequence, we obtain the following polynomial-time reduction.
@@
If $G$ is a direct product of simple groups, then $G$ $\Gamma(G)$ is uniquely determined by its commuting graph.
@@
Now suppose for a contradiction that $\Gamma(G)$ is the strong product of two nontrivial graphs with vertex sets $A$ and $B$ (i.e., with $|A|>1$ and $|B|>1$). Then we can identify $G$ with the Cartesian product $A\times B$. Suppose that $(a,b)$ is the identity of $G$. Then $(a,b)$ is joined to all other vertices in the commuting graph.
@@
Each case is contradictory; so our assumption that $\Gamma(G)$ is the strong product of two nontrivial graphs is false, and the theorem is proved.$\qed$
@@ 6 Commuting graphs of semidirect products
For example, consider the commuting graph of the dihedral group $D_{2n}$. Let $D_{2n}=\langle a,b\rangle$ where $a^{2}=1,b^{n}=1$ and $aba^{-1}=b^{-1}$. For $n$ odd, $1$ is the only dominating vertex, there is an $n$-clique corresponding to $\langle b\rangle$ and the $ab^{i}$ are pendant vertices. For $n$ even, it is a bit different with $b^{n/2}$ being the other dominating vertex.
More generally, consider the generalized dihedral group $\langle A,t\rangle$ defined earlier, where $A$ is abelian of twice odd order. Its commuting graph is isomorphic to that of the dihedral group of the same order. We conclude that we can recognise that a graph is the commuting graph of a generalized dihedral group, but cannot decide which one.
@@ 6.1 Commuting graphs of Frobenius groups
Proof 5
(a) By Thompson’s theorem, a Frobenius kernel is nilpotent, and nilpotent groups have a non-trivial centre.
(b) Let $H$ be a Frobenius complement. By [43, Theorem 6.3, Corollary 6.10], if $H$ has even order, then it has a unique element of order $2$, which lies in the centre; while, if $H$ has odd order, then it has a unique subgroup of order $p$ for every $p$ dividing $|H|$. In the latter case, choose $p$ to be the smallest prime divisor of $H$, and $P=\langle a\rangle$ the unique subgroup of order $p$. Then $P$ is a normal subgroup of $H$. For any $h\in H$, we have $a^{h}=a^{m}$ for some $m$ with $11\leq m\leq p-$ 1. If $m>1$, let $q$ be a prime divisor of $m$; then some power of $h$ has order $q$, a contradiction. So $a^{h}=a$ for all $h\in H$, whence $a\in Z(H)$$\qed$
@@
Let $G$ be a group of order $nk$, where $n,k>1$ and $\gcd(n,k)=1$, and let $\Gamma$ be the commuting graph of $G$. Then $G$ is a Frobenius group with Frobenius complement of order $k$ if and only if $\Gamma$ satisfies the following conditions:
@@
    there is a dominating vertex $v$;
@@
    $\Gamma\setminus\{v\}$ has a component of size $n-1$ and $n$ components of size $k-1$, and each component has a dominating vertex.
Proof 6
Suppose that $G$ is a Frobenius group with kernel $N$ of order $n$ and complement $H$ of order $k$. The identity is a dominating vertex. Also non-identity vertices in the kernel do not commute with non-identity vertices in a complement, and non-identity vertices in different complements do not commute with each other. So the components of $\Gamma\setminus\{1\}$ are as stated, and the lemmas above show that they have dominating vertices.
Conversely, suppose that the commuting graph $\Gamma$ of $G$ has properties (a) and (b). If $C$ is a component with a dominating vertex $c$, then the centralizer $C_{G}(c)$ is equal to $C\cup\{1\}$, which is thus a subgroup of $G$. Let $N$ be the subgroup containing the component of size $n-1$, and let $H_{1},\ldots,H_{n}$ be the subgroups containing the other components.
Since $\Gamma$ is invariant under automorphisms of $G$, we have an action of $G$ on the set $\Omega$ of components of $\Gamma\setminus\{v\}$ of size $k-1$ by conjugation. We show that this action satisfies the conditions for a Frobenius group given above.
Choose a prime $p$ dividing $k$. Since $p$ does not divide $n$, the subgroup $N$ cannot contain a Sylow $p$-subgroup; but each of $H_{1},\ldots,H_{n}$ contains such a subgroup. By the conjugacy part of Sylow’s theorem, $G$ acts transitively on $\Omega$.
A non-identity element of $G$, acting by conjugation, fixes itself, and so fixes the component containing it. We must show that it fixes no other component. Count fixed point of elements of $G$. The identity fixes $n$; non-identity elements of $N$ fix a positive number; and the remaining elements fix a non-negative number. So the sum of the fixed point numbers is at least $n+n(k-1)=nk$. But, by the Orbit-counting Lemma, this sum is equal to $|G|=nk$, since $G$ is transitive. So equality holds; non-identity elements of $N$ fix no point, and the remaining non-identity elements fix exactly one point each. So $G$ is a Frobenius group.$\qed$
@@
Above we assumed that our graph is a commuting graph. However, it is easy to construt a graph satisfying the conditions of Theorem 6.2 which is not a commuting graph. Can we find necessary and sufficient conditions for $\Gamma$ to be the commuting graph of a Frobenius group?
@@ 7 Recognizing commuting graphs of extraspecial groups
Our aim in this section is to design a polynomial-time algorithm that takes a simple undirected graph $X=(V,E)$, with $n=p^{2r+1}$ vertices, $r\geq 2$ and $p$ an odd prime, as input and determines if $X$ is the commuting graph of an extraspecial group $G$ of order $p^{2r+1}$. Moreover, the algorithm will also compute a certifying bijection $v\mapsto g_{v}$ labeling each vertex $v\in V$ by a unique group element $g_{v}\in G$ such that the commuting relations hold.
We recall that a $p$-group $G$ is extraspecial group $G$ if its centre $Z(G)$ is of order $p$ and coincides with its derived subgroup $G^{\prime}$ and the Frattini subgroup $\Phi(G)$ [43, Chapter 4, pp. 123]. Furthermore, it is known that there are exactly two non-isomorphic extraspecial groups of order $p^{2r+1}$, for each $r\geq 1$ and each prime $p$. These two groups, which we denote as $G_{1}$ and $G_{2}$, are described below. For a pair of elements $(x,y)$ in any group, let $[x,y]$ denote their commutator $x^{-1}y^{-1}xy$.
@@
    The first group $G_{1}$ is generated by $12r+$ 1 elements $G_{1}=\langle z,x_{i},y_{i},1\leq i\leq r\rangle$ such that $[x_{i},x_{j}]=1$, $[y_{i},y_{j}]=1$ for all $i,j$, and $[x_{i},y_{j}]=1$ for all $i\neq j$. $[x_{i},y_{i}]=z$ for all $i$, and $x_{i}^{p}=y_{i}^{p}=z^{p}=1$ for all $i$. The center $Z(G_{1})=\langle z\rangle$.
@@
    The second group $G_{2}$ also has $12r+$ 1 generators $G_{2}=\langle z,x_{i},y_{i},1\leq i\leq r\rangle$ such that $[x_{i},x_{j}]=1$, $[y_{i},y_{j}]=1$ for all $i,j$, and $[x_{i},y_{j}]=1$ for all $i\neq j$. And $[x_{i},y_{i}]=z$ for all $i$, and $x_{i}^{p}=1=z^{p}$, $y_{i}^{p}=z$ for all $i$. The center $Z(G_{2})=\langle z\rangle$.

### v4

@@ 3 The Recognition Problem
    For the general case, we obtain a quasipolynomial-time algorithm. The algorithm is essentially based on a simple $2^{O(\log^{3}n)}$ time algorithm that enumerates the multiplication tables of all distinct groups of order $n$ (where by distinct we mean mutually non-isomorphic groups of order $n$). This enumeration algorithm combined with Babai’s quasipolynomial time algorithm for graph isomorphism [4] yields a simple $2^{O(\log^{3}n)}$ time algorithm to check if a given $n$-vertex graph $X$ is the commuting graph of a group of order $n$: for each $n$-element group $G$ that is output by the enumeration algorithm we check if $\Gamma(G)$ is isomorphic to $X$ by using Babai’s isomorphism test [4].
@@ 4 Basic properties
Let $G$ be a finite group and $\Gamma(G)$ its commuting graph. A vertex subset $S$ is a maximal clique of $\Gamma(G)$ if and only if $S$ is a maximal abelian subgroup of $G$.
Proof 3
Suppose $H$ is a maximal abelian subgroup of $G$. Clearly, $H$ is a clique in $X$. If $x\notin H$ is adjacent to all of $H$ then $x$ commutes with all of $H$ implying that $\langle H\cup\{x\}\rangle$ is an abelian subgroup of $G$ strictly larger than $H$. Hence the clique induced by $H$ is maximal.
Conversely, suppose $S\subset G$ is a maximal clique in $\Gamma(G)$. Then any two elements $x,y\in S$ commute with each other. Hence, the subgroup $H=\langle S\rangle$ generated by $S$ is an abelian subgroup of $G$ which means that the vertex subset $H$ is a clique in $\Gamma(G)$ containing $S$. As $S$ is maximal, by assumption, it follows that $S=H$ which means $S$ is a maximal abelian subgroup. $\qed$
@@ 4.1 The vertex degrees of $\Gamma(G)$ and conjugacy classes of $G$
An $n$-vertex graph $X$ with fewer than $\frac{n(f(n)-1)}{2}$ edges cannot be the commuting graph of a group, where $f(n)$ is the minimum number of conjugacy classes that a group of order $n$ can have.
At the other extreme, we can rule out very dense incomplete graphs by the $5/8$-theorem [43] for finite groups. More precisely, for any finite non-abelian group $G$ with $n$ elements:
@@ 4.2 The commuting graph and maximal abelian subgroups
For a finite group $G$, the number of maximal abelian subgroups is bounded by (${|G|}\choose{\log|G|}$ because every subgroup of $G$ has a generating set of size bounded by $\log|G|$.
@@ 5 Commuting graphs of product groups
[44, Chapter 4, Page 35] Let $X=(V,E)$ and $X^{\prime}=(V^{\prime},E^{\prime})$ be simple undirected graphs. Their strong product, denoted $X\boxtimes X^{\prime}$, is a simple undirected graph with the cartesian product $V\times V^{\prime}$ as its vertex set and edges defined as follows: distinct pairs $(u,u^{\prime})$ and $(v,v^{\prime})$ are adjacent if and only if one of the following holds:
@@
The strong product is both commutative and associative [44, Chapter 4], in the sense that $X\boxtimes Y\cong Y\boxtimes X$ and $(X\boxtimes Y)\boxtimes Z\cong X\boxtimes(Y\boxtimes Z)$.
By a non-trivial graph we mean a graph with more than one vertex. Given a graph $X$, a factorization $X\cong Y\boxtimes Z$ is non-trivial if both $Y$ and $Z$ are non-trivial. A graph $X$ is said to be a prime graph if it does not have a non-trivial factorization. (We warn the reader that this term is also used for graphs defined by arithmetic parameters of a finite group; see for example [55].)
It turns out that any connected graph $X$ has a unique factorization (up to isomorphism) into a strong product of prime graphs [44, Theorem 7.14, Pg 85]. Moreover, this prime factorization can be computed in polynomial time [35]. As a consequence, we obtain the following polynomial-time reduction.
@@
If $G$ is a direct product of simple groups, then $G$ is uniquely determined by its commuting graph $\Gamma(G)$.
@@
Now suppose for a contradiction that $\Gamma(G)$ is the strong product of two non-trivial graphs with vertex sets $A$ and $B$ (i.e., with $|A|>1$ and $|B|>1$). Then we can identify $G$, as the vertex set of $\Gamma(G)$, with the Cartesian product $A\times B$. Suppose that $(a,b)$ is the identity of $G$. Then $(a,b)$ is joined to all other vertices in the commuting graph.
@@
Each case is contradictory; so our assumption that $\Gamma(G)$ is the strong product of two non-trivial graphs is false, and the theorem is proved.$\qed$
@@ 6 Commuting graphs of semidirect products
For example, consider the commuting graph of the dihedral group $D_{2n}$. Let $D_{2n}=\langle a,b\rangle$ where $a^{n}=b^{2}=1$ and $bab=a^{-1}$. For $n$ odd, $1$ is the only dominating vertex, there is an $n$-clique corresponding to $\langle a\rangle$ and the $ba^{i}$ correspond to pendant vertices. Thus, the commuting graph of $D_{2n}$ is easily recognizable. For $n$ even, it is a bit different with $b^{n/2}$ being the other dominating vertex.
More generally, consider the generalized dihedral group $\langle A,t\rangle$ defined earlier, where $A$ is abelian of order $2n$, $n$ odd. Its commuting graph is isomorphic to that of the dihedral group $D_{2n}$ of the same order. Hence, as observed above, we can recognize that a graph is the commuting graph of a generalized dihedral group, but cannot decide which one.
@@ 6.1 Commuting graphs of Frobenius groups
Proof 7
(a) By Thompson’s theorem [76, Theorem 1], a Frobenius kernel is nilpotent, and nilpotent groups have a non-trivial centre.
(b) Let $H$ be a Frobenius complement. By [45, Theorem 6.3, Corollary 6.10], if $H$ has even order, then it has a unique element of order $2$, which lies in the centre; while, if $H$ has odd order, then it has a unique subgroup of order $p$ for every $p$ dividing $|H|$. In the latter case, choose $p$ to be the smallest prime divisor of $H$, and $P=\langle a\rangle$ the unique subgroup of order $p$. Then $P$ is a normal subgroup of $H$. For any $h\in H$, we have $a^{h}=a^{m}$ for some $m$ with $11\leq m\leq p-$ 1. If $m>1$, let $q$ be a prime divisor of $m$; then some power of $h$ has order $q$, a contradiction. So $a^{h}=a$ for all $h\in H$, whence $a\in Z(H)$$\qed$
@@
Let $G$ be a group of order $nk$, where $n>1$, $k>1$, and $n$ and $k$ are relatively prime. Then $G$ is a Frobenius group with Frobenius complement of order $k$ if and only if its commuting graph $\Gamma(G)$ satisfies the following conditions:
@@
    there is a unique dominating vertex $1$ in $\Gamma(G)$, where $1$ corresponds to the identity element of $G$.
@@
    $\Gamma(G)\setminus\{1\}$ has a component of size $n-1$ and $n$ components of size $k-1$, and each component has a dominating vertex.
Proof 8
Suppose that $G$ is a Frobenius group with kernel $N$ of order $n$ and complement $H$ of order $k$. The identity is a dominating vertex. Also non-identity vertices in the kernel do not commute with non-identity vertices in a complement, and non-identity vertices in different complements do not commute with each other. Thus the components of $\Gamma(G)\setminus\{1\}$ are as stated, and Lemma 6.1 shows that they have dominating vertices.
Conversely, suppose that the commuting graph $\Gamma(G)$ has properties (a) and (b). If $C$ is a component with a dominating vertex $c$, then the centralizer $C_{G}(c)$ is equal to $C\cup\{1\}$, which is thus a subgroup of $G$. Let $N$ be the subgroup containing the component of size $n-1$, and let $H_{1},\ldots,H_{n}$ be the subgroups containing the other components. Note that $|N|=n$ and $|H_{i}|=k,1\leq i\leq n$.
Since $\Gamma(G)$ is invariant under automorphisms of $G$, we have an action of $G$ on the set $\Omega$ of components of $\Gamma(G)\setminus\{1\}$ of size $k-1$ by conjugation. We show that this action satisfies the conditions for a Frobenius group given above.
Choose a prime $p$ dividing $k$. Since $n$ and $k$ are relatively prime, $p$ does not divide $n$. So the subgroup $N$ cannot contain a Sylow $p$-subgroup; but each of $H_{1},\ldots,H_{n}$ contains such a subgroup. By the conjugacy part of Sylow’s theorem, $G$ acts transitively on $\Omega$.
For $1\leq i\leq n$, each element of $H_{i}\setminus\{1\}$ acting by conjugation, fixes itself and so fixes the component induced by $H_{i}\setminus\{1\}$. That is, a non-identity element of $G\setminus N$, acting by conjugation, fixes itself, and so fixes the component containing it. We must show that it fixes no other component. To that end, we count the fixed points of elements of $G$ acting on $\Omega$. The identity fixes $n$; each non-identity element of $N$ fixes some non-negative number; and each of the remaining elements $G\setminus N$ fix a positive number of elements. So the sum of the fixed point numbers is at least $n+n(k-1)=nk$. But, by the Orbit-counting Lemma, this sum is equal to $|G|=nk$, since $G$ is transitive. Consequently, non-identity elements of $N$ do not fix any point of $\Omega$, and the remaining non-identity elements fix exactly one point each. So $G$ is a Frobenius group.$\qed$
@@
Above we assumed that our graph is a commuting graph. However, it is easy to construct a graph satisfying the conditions of Theorem 6.2 which is not a commuting graph. Can we find necessary and sufficient conditions for $\Gamma$ to be the commuting graph of a Frobenius group?
@@ 7 Recognizing commuting graphs of extraspecial groups
In this section we design a polynomial-time algorithm that takes a simple undirected graph $X=(V,E)$, with $n=p^{2r+1}$ vertices, $r\geq 2$ and $p$ an odd prime, as input and determines if $X$ is the commuting graph of an extraspecial group $G$ of order $p^{2r+1}$ (see Section 2) for definition). The algorithm will also compute a certifying bijection $v\mapsto g_{v}$ labeling each vertex $v\in V$ by a unique group element $g_{v}\in G$ such that the commuting relations hold.
There are exactly two non-isomorphic extraspecial groups of order $p^{2r+1}$, for each $r\geq 1$ and each prime $p$. These two groups, which we denote as $G_{1}$ and $G_{2}$, are described below.
@@
    The first group $G_{1}$ is generated by $12r+$ 1 elements $G_{1}=\langle z,x_{i},y_{i},1\leq i\leq r\rangle$ such that $[x_{i},x_{j}]=1$, $[y_{i},y_{j}]=1$ for all $i,j$, and $[x_{i},y_{j}]=1$ for all $i\neq j$, $[x_{i},y_{i}]=z$ for all $i$, and $x_{i}^{p}=y_{i}^{p}=z^{p}=1$ for all $i$. The centre $Z(G_{1})=\langle z\rangle$.
@@
    The second group $G_{2}$ also has $12r+$ 1 generators $G_{2}=\langle z,x_{i},y_{i},1\leq i\leq r\rangle$ such that $[x_{i},x_{j}]=1$, $[y_{i},y_{j}]=1$ for all $i,j$, and $[x_{i},y_{j}]=1$ for all $i\neq j$. And $[x_{i},y_{i}]=z$ for all $i$, and $x_{i}^{p}=1=z^{p}$, $y_{i}^{p}=z$ for all $i$. The centre $Z(G_{2})=\langle z\rangle$.